The sum of the present ages of Arie and her mother is 64 years. In eight years, the sum of their ages will be three-fifths as old as her mother at that time. find the present age of Arie and her mother. (please use only one variable and thanks in advance)
Found 2 solutions by VFBundy, greenestamps:
Answer by VFBundy(438)

The sum of the present ages of Arie and her mother is 64 years:

Arie's age = A
Mother's age = 64-A

In eight years, the sum of their ages will be three-fifths as old as her mother at that time:

3/5*((A+8) + (64-A+8)) = 64-A+8

Simplify:

3/5*(80) = 72-A

48 = 72-A

-24 = -A

A = 24

Arie's age = A = 24
Mother's age = 64-A = 40

Answer by greenestamps(13200)



Very badly worded problem; and, exactly as stated, it is nonsense.

"In eight years, the sum of their ages will be three-fifths as old as her mother at that time."

(1) "the sum of their ages will be three-fifths as old as..." says that the sum has an age. But a sum is a number; it doesn't have an age. The sum of the two ages can't be "3/5 as old as" one of the ages.

(2) If the sum of two ages is 3/5 of one of the two ages, then the second age is a negative number -- which is clearly nonsense.